OpenStudy (anonymous):

solve. -112/3 x (-4 1/5) = ? A. 44 2/15 B. -44 2/15 C. 49 D. -49

OpenStudy (anonymous):

Rule for Multiplying Fractions:( a/b) x (c/d) = ac/bd

OpenStudy (anonymous):

Step 1: convert all numbers to form (a/b) and (c/d)

OpenStudy (anonymous):

Applying Step 1, we get: (a/b) = (-112/3) (c/d) = (-21/5)

OpenStudy (anonymous):

Step 2: Apply Rule for Multiplying Fractions (above) Step 3: Reduce/Simplify the Answer

OpenStudy (anonymous):

For Step 2, write your answer as (-112)(-21)/(3)(5) For Step 3, simplify (-21/3) = -7 then the answer is just (-112)(-7))/(5)!